# Adversarial Inception for Bounded Backdoor Poisoning in Deep Reinforcement Learning

### v1

@@ 5 Adversarial Inception Algorithm
1:Policy $\pi$, Replay Memory $\mathcal{D}$, max episodes $N$, Lower Bound $\hat{L}$, Upper Bound $\hat{U}$
2:training algorithm $\mathcal{L}$, benign MDP $M=(S,A,R,T,\gamma)$, poisoning rate $\beta$, trigger $\delta$
3:for $i\leftarrow 1,N$ do
4:     Victim samples trajectory $H=\{(s,a,r)_{t}\}_{t=1}^{\mu}$ of size $\mu$ from $M$ given policy $\pi$
5:     Update $\hat{L}\leftarrow\min[\hat{L},\min[r_{t}]]$, $\hat{U}\leftarrow\max[\hat{U},\max[r_{t}]]$
6:     Select $H^{\prime}\subset H$ using metric $\mathcal{F}_{\hat{Q}}(s_{t},a_{t})$ s.t. $|H^{\prime}|=\lfloor\beta\cdot|H|\rfloor$
7:     for all $(s,a,r)_{t}\in H^{\prime}$ do
8:         $s_{t}\leftarrow\delta(s_{t})$, $r_{old}\leftarrow r_{t}$
9:         $a_{t}\leftarrow a^{+}$ if $\mathcal{F}_{\hat{Q}}(s_{t},a_{t})>0$
10:         $r_{t}\leftarrow U$ if $a_{t}=a^{+}$ else $L$
11:         $r_{t-1}\leftarrow\text{clip}(r_{t-1}-\gamma(r_{t}-r_{old}),L,U)$
12:     Victim stores perturbed $H$ in $\mathcal{D}$ then updates $\pi$ with $\mathcal{L}$ given $\mathcal{D}$,
13:     Update $\hat{Q}$ for metric $\mathcal{F}_{\hat{Q}}$ given $\mathcal{D}$ using DQN
In Algorithm 1 we present a generalized framework for inception attacks against DRL with the aim of replicating the adversarial MDP $M^{\prime}$ specified in Section 4. In $M^{\prime}$ we use $\phi$ to force the MDP to transition with respect to optimal actions when the agent chooses target action $a^{+}$, however this isn’t possible under our threat model. The adversary does not have direct access to $Q_{\pi}^{M^{\prime}}(s)$ nor can they change the agent’s actions during an episode $H=\{(s,a,r)_{t}\}$. Thus we take a more indirect approach in steps 6 and 7 – incepting false values in the agent replay memory $\mathcal{D}$ so they think they took action $a^{+}$ in poisoned state $\delta(s_{t})$ when, in reality, they chose and transitioned with respect to some high value action $a_{t}$ in benign state $s_{t}$. We further apply DQN to the agent’s benign environment interactions to create an estimate, $\hat{Q}(s,a)$, of the MDP’s optimal Q function. With this we can create a metric $\mathcal{F}_{\hat{Q}}$, like the one defined in Equation 15 for Q-Incept, to approximate the relative optimality of each action.
@@
This allows us to approximate $\phi$ by finding time steps in which the agent took near optimal actions in step 4. Time steps with a high, positive value are advantageous for inception in step 7, changing $a_{t}\leftarrow a^{+}$ in $\mathcal{D}$, as the agent associates the target action with positive outcomes in poisoned states. Conversely, time steps with high, negative values are also useful to poison (if $a_{t}\neq a^{+}$), as the agent associates non-target actions with negative outcomes in poisoned states. In Q-Incept we use the absolute value of $\mathcal{F}_{\hat{Q}}(s,a)$ as softmax logits to weigh how we sample $H^{\prime}\subseteq H$ in step 4. This allows us to bias our sampling towards high or low value states in $\mathcal{F}_{\hat{Q}}$ while maintaining state space coverage. In steps 8-9 we opt to implement a slightly stronger version of $\tau$ which perturbs the agent’s rewards to $U$ or $L$ if $a_{t}=a^{+}$ or $a_{t}\neq a^{+}$ respectively. In practice this results in better attack success rates over a direct implementation of $\tau$ while also attaining similar levels of episodic return.
@@ 6 Experimental Results
Here we evaluate Q-incept against TrojDRL and SleeperNets, representing the state of the art in forced action manipulation and unbounded reward poisoning attacks respectively. All attacks are evaluated under bounded reward poisoning, restricting each attack to reward perturbations that lie within the observed min and max of the agent’s benign rewards. We evaluate these attacks using cleanrl’s implementation of PPO (Huang et al., 2022) on 4 environments: CAGE-2 (Kiely et al., 2023), Highway Merge (Leurent, 2018), Q*Bert (Brockman et al., 2016), and Safety Car (Ji et al., 2023) which are cyber-network defending, simulated self-driving, Atari game playing, and robotic navigating tasks respectively. This diverse set of domains allows us to verify the effectiveness of Q-Incept across tasks with little overlap. All attacks are evaluated under the same poisoning budget unique to each environment and averaged over 5 seeds. In each environment we compare against a baseline agent trained without poisoning to compute our attack stealth metric, Benign Return Ratio (BRR) (Rathbun et al., 2024), as the ratio between the poisoned and unpoisoned agents’ episodic returns. Further experimental details and results are given in the appendix.
Empirical Comparison of Backdoor Attacks Under Bounded Rewards
Environment CAGE-2 ($\beta=1\%$) Highway Merge ($\beta=10\%$) Qbert ($\beta=0.03\%$) Safety Car ($\beta=0.1\%$)
Mertric ASR $\sigma$ ASR $\sigma$ ASR $\sigma$ ASR $\sigma$
Q-Incept 93.21% 15.13% 61.60% 23.29% 100 % 0.00% 100 % 0.00%
SleeperNets 0.06% 0.12% 1.50% 0.53% 55.61% 39.35% 83.95% 29.12%
TrojDRL 5.64% 7.73% 1.20% 0.67% 22.51% 20.77% 53.35% 9.51%
Mertric BRR $\sigma$ BRR $\sigma$ BRR $\sigma$ BRR $\sigma$
Q-Incept 99.29% 17.27% 97.97% 1.00% 100% 8.07% 80.48% 3.34%
SleeperNets 88.92% 18.86% 99.89% 0.08% 98.48% 11.59% 80.94% 17.39%
TrojDRL 82.43% 19.29% 99.65% 0.20% 100% 5.32% 86.37% 9.65%
Table 5: Comparison between Q-Incept, SleeperNets, and TrojDRL with bounded rewards against agents training on CAGE-2, Highway Merge, Q*Bert, and Safety Car. Attacks are evaluated on Attack Success Rate (ASR) and Benign Return Ratio (BRR). Attacks with the highest average BRR or ASR on each environment are printed in bold. Standard deviations $\sigma$ are given next to each result.
In Table 5 we present our results. Across all four environments Q-Incept out-performs both SleeperNets and TrojDRL in terms of average attack success rate (ASR) over 5 runs. Of note is Cage-2 were Q-Incept achieves 93.21% ASR, while TrojDRL and SleeperNets both fail to achieve above 6%. Highway Merge was the most difficult environment for all attacks, likely due to its high penalty for crashing and short training time, however Q-Incept was that only method that scaled with $\beta$, achieving an ASR of 61.6% while the other attacks fall below 2% even at $\beta=10\%$. These cases highlight how, under bounded reward poisoning, prior attacks like SleeperNets and TrojDRL can become completely ineffective. Even in cases where SleeperNets and TrojDRL achieve decent levels of attack success, like on Q*Bert and Safety Car, Q-Incept still out performs them - consistently attaining an ASR of 100% on both environments. SleeperNets was able to achieve near 95% ASR on Q*Bert across all but one seed, making it effective but less reliable than Q-Incept. We also see that our attack stealth metric, BRR, is relatively consistent across attacks - with each achieving nearly 100% BRR across all environments excluding Safety Car. This demonstrates how Q-Incept achieves both our objectives of attack success and attack stealth during real-time attacks - verifying our theoretical analysis of adversarial inception.
6.1 Ablations
Here we perform some ablations to further study two observations we made. First, on CAGE-2 in particular Q-Incept was able to greatly out-perform the other two attacks not leveraging adversarial inception. To verify that adversarial inception is the main contributor for this success we compare against Q-Vanilla, which simply skips step 7 in the Q-Incept algorithm, resulting in no inception. In Figure 4 we see that this modification results in a significant drop from 93.21% ASR to under 5%, indicating that adversarial inception is crucial for the success of Q-Incept.
@@ 6.1 Ablations
Figure 4: Comparison between Q-Incept with and without adversarial inception on the CAGE-2 environment. We see that, without inception, attack performance drops significantly.
Next, we noticed that Highway Merge was the only environment on which Q-Incept was unable to attain an average ASR above 90%, leading us to question if our Q-function based approach was incorrect or if our online DQN approximation $\hat{Q}$ wasn’t converging quickly enough. To test this we devised Oracle-Incept – which uses an oracle Q-function pre-trained with DQN until convergence – as a hypothetical, stronger attack by an adversary with direct access to the benign MDP. In Figure 5 we can see that Oracle-Incept improves greatly over Q-Incept, reaching an average ASR of 93.38%. This indicates that better Q-function approximations lead to better performance - validating that both our chosen metric and attack approach scale properly with the accuracy of $\hat{Q}$. Thus, adversaries capable of using Q-function estimations with faster convergence can expect greater attack success.
Figure 5: Comparison between Q-Incept, Oracle-Incept, TrojDRL, and SleeperNets on Highway Merge. We can see that the Oracle-Incept shows significant improvements in ASR.
7 Conclusion
@@
In this paper we provide multiple contributions towards a deeper understanding of backdoor poisoning attacks against DRL algorithms. We demonstrate how prior attacks are detectable and how attempts to evade detection via reward clipping results in attack failure. We then propose Adversarial Inception as a novel framework for backdoor poisoning attacks against DRL under bounded reward perturbation constraints. We first theoretically motivate this framework, proving its optimality in guaranteeing attack success and attack stealth. We then develop an approximate adversarial inception attack, Q-Incept, which achieves state-of-the-art performance for different environments from different domains, while remaining stealthy. This novel threat necessitates future research into techniques for detecting and mitigating adversarial inception attacks along with further explorations into the capabilities of increasingly realistic and stealthy threat models and attack formulations.

### v3

@@ 5 Adversarial Inception Algorithm
Policy $\pi$, Replay Memory $\mathcal{D}$, max episodes $N$, Empirical Lower Bound $\hat{L}$, Empirical Upper Bound $\hat{U}$algorithm $\mathcal{L}$, MDP $M$, poisoning rate $\beta$, trigger $\delta$$i\leftarrow 1,N$Victim samples trajectory $H=\{(s,a,r)_{t}\}_{t=1}^{\mu}$of size $\mu$from $M$given policy $\pi$$\hat{L}\leftarrow\min[\hat{L},\min[r_{t}]]$, $\hat{U}\leftarrow\max[\hat{U},\max[r_{t}]]$Select $H^{\prime}\subset H$of size $\lfloor\beta\cdot|H|\rfloor$using $\mathcal{F}_{\hat{Q}}(s_{t},a_{t})$$(s,a,r)_{t}\in H^{\prime}$$s_{t}\leftarrow\delta(s_{t})$, $r_{old}\leftarrow r_{t}$$a_{t}\leftarrow a^{+}$if $\mathcal{F}_{\hat{Q}}(s_{t},a_{t})>0$$r_{t}\leftarrow\hat{U}$if $a_{t}=a^{+}$else $\hat{L}$$r_{t-1}\leftarrow\min[\max[r_{t-1}-\gamma(r_{t}-r_{old}),\hat{L}],\hat{U}]$Victim stores $H$in $\mathcal{D}$, updates $\pi$with $\mathcal{L}$given $\mathcal{D}$Update $\hat{Q}$for metric $\mathcal{F}_{\hat{Q}}$given $\mathcal{D}$using DQN
[…]
In Algorithm 1, we present a framework for inception attacks against DRL with the aim of replicating the adversarial MDP $M^{\prime}$ from Section 4. In $M^{\prime}$ we use $\phi$ to force the MDP to transition with respect to optimal actions when the agent chooses target action $a^{+}$, but this is not possible under our threat model. The adversary does not have direct access to $Q_{\pi}^{M^{\prime}}(s)$ nor can they change the agent’s actions during an episode $H=\{(s,a,r)_{t}\}$. Thus we take a more indirect approach in steps 6 and 7 – incepting false values in the agent replay memory $\mathcal{D}$ so they think they took action $a^{+}$ in poisoned state $\delta(s_{t})$ when, in reality, they chose and transitioned with respect to some high value action $a_{t}$ in benign state $s_{t}$. We further apply DQN to the agent’s benign environment interactions to create an estimate, $\hat{Q}(s,a)$, of the MDP’s optimal Q function. With this, we can create a metric $\mathcal{F}_{\hat{Q}}$, like the one defined in Equation 13 for Q-Incept, to approximate the relative optimality of each action.
@@
where we approximate the agent’s current policy $\pi$ using the maximum entropy, softmax distribution (Ziebart et al., 2008) $\pi(s^{\prime},\cdot)\approx\text{softmax}[\hat{Q}(s^{\prime},\cdot)]$. This construction allows us to approximate $\phi$ by finding time steps in which the agent took near optimal actions in step 4. Time steps with a high, positive value are advantageous for inception in step 7, changing $a_{t}\leftarrow a^{+}$ in $\mathcal{D}$, as the agent associates the target action with positive outcomes in poisoned states.
Environment Qbert Frogger Pacman Breakout CAGE-2 Highway Merge Safety Goal
Target Action Move Right Move Down No-Op No-Op No-Op Merge Right Accelerate
$\beta$ 0.3% 0.3% 0.3% 0.1% 1% 10% 0.1%
Metric ASR $\sigma$ ASR $\sigma$ ASR $\sigma$ ASR $\sigma$ ASR $\sigma$ ASR $\sigma$ ASR $\sigma$
Q-Incept 100% 0.0% 100% 0.0% 100.0% 0.0% 100% 0.0% 93.21% 15.13% 61.60% 23.29% 100% 0.00%
SleeperNets 55.6% 39.3% 0.00% 0.00% 11.3% 6.9% 100% 0.0% 0.06% 0.12% 1.50% 0.53% 86.96% 29.12%
TrojDRL 22.5% 20.7% 4.42% 9.88% 13.5% 4.5% 99.3% 0.4\$ 5.64% 7.73% 1.20% 0.67% 54.04% 2.85%
Metric BR $\sigma$ BR $\sigma$ BR $\sigma$ BR $\sigma$ BR $\sigma$ BR $\sigma$ BR $\sigma$
Q-Incept 18,381 882 392.3 37.4 583.0 116.2 460.8 20.3 -45.49 8.04 14.85 0.15 12.00 0.16
SleeperNets 16,840 1,982 366.5 50.9 665.5 281.7 443.4 13.8 -50.18 8.52 15.14 0.01 9.23 1.98
TrojDRL 17,617 909 366.3 33.9 670.7 189.4 455.3 16.8 -53.11 8.71 15.11 0.03 10.86 1.13
$\beta$ 0.1% 0.1% 0.1% 0.05% 0.5% 7.5% 0.05%
Metric ASR $\sigma$ ASR $\sigma$ ASR $\sigma$ ASR $\sigma$ ASR $\sigma$ ASR $\sigma$ ASR $\sigma$
Q-Incept 100% 0.00% 89.5% 1.16% 60.1% 46.1% 100% 0.0% 30.61% 15.01% 53.03% 25.56% 100% 0.00%
SleeperNets 19.98% 4.39% 45.92% 9.17% 42.7% 41.4% 100% 0.0% 0.00% 0.00% 1.47% 0.84% 83.95% 13.45%
TrojDRL 15.38% 3.05% 44.00% 10.63% 48.9% 30.3% 99.1% 0.5% 0.00% 0.00% 3.27% 3.56% 53.35% 9.51%
Metric BR $\sigma$ BR $\sigma$ BR $\sigma$ BR $\sigma$ BR $\sigma$ BR $\sigma$ BR $\sigma$
Q-Incept 17,749 1,380 437.9 10.4 457.1 87.3 456.1 19.7 -44.19 16.68 14.90 0.13 11.52 0.40
SleeperNets 17,320 1,481 412.8 11.8 584.9 89.0 456.5 15.5 -57.82 8.38 15.15 0.01 10.4 0.7
TrojDRL 17,715 1340 357.0 52.1 712.0 191.6 443.5 25.9 -39.28 9.60 15.1 0.02 9.8 1.1
Metric BR $\sigma$ BR $\sigma$ BR $\sigma$ BR $\sigma$ BR $\sigma$ BR $\sigma$ BR $\sigma$
No Poisoning 17,322 1,773 380.4 89.1 628.7 236.5 476.5 9.4 -45.17 15.65 15.17 0.01 12.03 0.16
Table 3: Comparison between Q-Incept, SleeperNets, and TrojDRL with constrained rewards against agents training on our seven environments at different $\beta$ values. Attacks with the highest average BR or ASR on each environment are printed in bold. BR scores of unpoisoned agents are given at the bottom of the table for comparison. Standard deviations $\sigma$ are given next to each result.
Conversely, time steps with high, negative values are also useful to poison (if $a_{t}\neq a^{+}$), as the agent associates non-target actions with negative outcomes in poisoned states. In Q-Incept we use the absolute value of $\mathcal{F}_{\hat{Q}}(s,a)$ as softmax logits to weigh how we sample $H^{\prime}\subseteq H$ in step 4. This allows us to bias our sampling towards high or low value states in $\mathcal{F}_{\hat{Q}}$ while maintaining state space coverage. In steps 8-9 we opt to implement a slightly stronger version of $\tau$ which perturbs the agent’s rewards to $\hat{U}$ or $\hat{L}$ if $a_{t}=a^{+}$ or $a_{t}\neq a^{+}$, respectively. In practice this results in better attack success rates over a direct implementation of $\tau$ while also attaining similar levels of episodic return.
@@ 6 Experimental Results
Here, we evaluate Q-Incept against TrojDRL and SleeperNets, representing the state of the art in forced action manipulation and dynamic reward poisoning attacks. We perform our evaluation in terms of Attack Success Rate (ASR) and Benign Return (BR), relating to our objectives of attack success and attack stealth, respectively, defined below:
$\begin{array}[]{c}\textbf{ASR}(\pi^{+}|\delta)\doteq\mathbb{E}_{s\in S}[\pi^{+}(\delta(s))]\\ \textbf{BR}(\pi^{+}|M)\doteq\mathbb{E}_{s_{0}\sim M}[V_{\pi^{+}}^{M}(s_{0})]\end{array}$ (14)
@@
where $s_{0}$ is a (potentially random) initial state given by $M$ and $\pi^{+}$ is the poisoned policy we are evaluating. Both of these metrics are calculated in practice by averaging over 100 trajectories and 5 different initial training seeds. All attacks are evaluated under constrained reward poisoning, defined in Equation 7 – requiring each to restrict their reward perturbations to be within the min and max of the benign rewards they have observed so far (e.g., lines 3 and 9 in Algorithm 1). We evaluate these attacks using cleanrl’s implementation of PPO (Huang et al., 2022) on 7 environments. Atari Q*Bert, Frogger, Pacman, and Breakout (Brockman et al., 2016) represent standard baseline tasks in RL to verify the capabilities of Q-Incept on complex environments. Additionally, CAGE Challenge 2 (Kiely et al., 2023), Highway Merge (Leurent, 2018), and Safety Car (Ji et al., 2023) extend the diversity of our analysis to other domains spanning cyber network defending, simplified self driving, and safety-aware robotic navigation tasks, respectively. This allows us to verify the effectiveness of Q-Incept across multiple task domains which share little overlap. Further experimental details and results are given in Appendix A.2 and Appendix A.4, respectively.
In Table 3, we present the performance of our attack on each environment across two different poisoning rates $\beta$ each. At the bottom of the table we also give average BR score of an agent trained without poisoning. Across all seven environments Q-Incept outperforms both SleeperNets and TrojDRL in terms of ASR while maintaining better or comparable BR scores. We also see that Q-Incept is the only method which consistently scales in terms of ASR as the poisoning budget $\beta$ increases – with SleeperNets and TrojDRL sometimes dropping in performance under a larger $\beta$ in cases such as Atari Frogger and Pacman.
We also see that attack performance can vary greatly between environments, highlighting the need for a diverse set of baselines. Atari Breakout, a baseline also studied by Rathbun et al. (2024) and Kiourti et al. (2019), seems to be the easiest environment to poison by far. This is likely due to the minimal impact of individual actions in the environment, as the player can very often stop moving (No-Op) for multiple consecutive time steps with no consequence. Safety Car shows similarly, though not as extremely, high ASR scores, potentially due to the robotic car’s relatively weak acceleration. Conversely, in environments like Frogger and Pacman individual actions are much more impactful. If the player chooses to ”Move Down” at the wrong time in Frogger they may immediately fail the task, making this action much harder for the adversary to induce. Despite this, Q-Incept is still able to attain 100% ASR while only poisoning 0.3% of time steps in both Frogger and Pacman.
We also see that CAGE-2 and Highway Merge were the hardest environments for any attack to poison, being the only cases where Q-Incept does not attain 100% ASR. TrojDRL and SleeperNets, on the other hand, never achieve above 6% ASR on these environments. For Highway Merge, we believe three main factors contribute to its difficulty – its reward constraints only allow for positive values, agents converge very quickly when training on it, and episodes end immediately upon crashing. Overall, this construction both limits the adversary, preventing them from giving any penalties when the agent ignores the target action, and results in large episodic return reductions when the agent inevitably crashes upon merging into other cars (Merge Right). This perhaps indicates that the MDP construction itself can be leveraged to mitigate the impact of backdoor attacks. For CAGE-2 the reasons are less clear, though we suspect its relatively large action space and the universal sub-optimality of the target action (No-Op) may be contributing factors.
6.1 Verifying the Importance of Inception
@@
Figure 3: Comparison between Q-Incept with and without adversarial inception on the CAGE-2 environment at $\beta=1\%$. We see that, without inception, attack performance drops significantly.
Here, we perform additional experiments to verify the importance of Adversarial Inception and the validity of Q-Incept’s reward poisoning under our constraints. First, in Figure 3 we compare Q-Incept both with and without inception (step 7 of Algorithm 1) on CAGE-2, referring to this naive version as “Q-Vanilla”. Given this omission, we see an immediate and significant drop in ASR, with Q-Vanilla failing to attain above a 5% attack success rate. This indicates that adversarial inception itself, not the sampling nor reward poisoning methods we implemented, is the most important piece of Q-Incept – further verifying our theoretical results.
Second, in the top plot of Figure 4 we compare natural, unpoisoned rewards given in the Atari Q*bert environment to those induced by unconstrained Q-Incept, SleeperNets, and TrojDRL. Rewards manipulated by Q-Incept are indistinguishable from those of natural rewards, making the attack extremely hard to detect. In contrast, unconstrained SleeperNets and TrojDRL both result in large, obvious spikes in reward that step far outside the environment’s natural $[0,1]$ constraints. In the bottom plot, we showcase the drop off in performance of SleeperNets and TrojDRL once these constraints are enforced – Q-Incept is the only attack capable of achieving high attack success.
Figure 4: (Top) Rewards induced by SleeperNets, TrojDRL, and Q-Incept compared to benign behavior on the Atari Q*bert Environment. (Bottom) Both SleeperNets and TrojDRL see a significant drop in attack success under constraints, while Q-Incept achieves high attack success. Attacks are given a budget $\beta=0.3\%$.
6.2 Evading Trigger-Based Universal Defenses
Many defenses against backdoor attacks in DRL, such as Bird (Chen et al., 2024) and State Sanitization (Bharti et al., 2022), claim universal defense capabilities by detecting or sanitizing the adversarial trigger directly at test time. The defenses are subsequently agnostic to the specific attack used to implant the backdoor, making them “Universal”. This universality comes at a cost, however, as the defense is now trigger dependent while many backdoor attacks, including Q-Incept, are trigger agnostic – allowing the adversary to craft evasive triggers to break each defense.
Figure 5: Example poisoned state in the Atari Q*bert environment before (left) and after (right) state sanitization. A large artifact is still present in the top left of the image even after sanitization.
For instance, the state sanitization defense proposed by Bharti et al. (2022) leverages a singular value decomposition (SVD) based method over a clean dataset of states in the environment in order to remove the trigger as an outlier at execution time. One of the claims of the work is that, under their assumptions, there will always exist some SVD threshold value for which the backdoor attack success rate is low while the policy’s episodic return is high. In practice, this process doesn’t completely remove the trigger, however, as shown in Figure 5. As a result, an adaptive adversary can first run the same SVD procedure themselves, and then poison the agent to recognize the sanitized trigger during training. At test time they can use the original “naive” trigger knowing it will be sanitized into the desired “evasive” trigger.
Figure 6: Comparing the performance of Q-Incept against the state sanitization defense with evasive and naive triggers. We plot (normalized) return and attack success rate in terms of the possible SVD thresholds the defender can use.
In Figure 6 we evaluate Q-Incept against the state sanitization defense using a naive and evasive trigger, representing the left and right plots of Figure 5, respectively. Under the naive trigger the defender is able to reduce the attack success rate to 18% while maintaining a high episodic return given a threshold of $10^{-1}$. Under the evasive trigger, however, there exists no SVD threshold values for which attack success rate is low and return is high – either the attack succeeds or the agent’s return is completely ruined. Therefore the defense is broken.
7 Conclusion and Discussion
@@
In this paper, we provide multiple contributions towards a deeper understanding of backdoor poisoning attacks against DRL algorithms. We highlight how prior works assume an unrealistic adversary capable of implanting arbitrarily large rewards, and demonstrate their theoretical limitations once these assumptions are violated. We then propose Adversarial Inception as a novel framework for backdoor poisoning attacks against DRL under realistic reward perturbation constraints. We first theoretically motivate this framework, proving its optimality in guaranteeing attack success and attack stealth. We then develop a practical adversarial inception attack, Q-Incept, which achieves state-of-the-art performance on multiple environments from different domains, while remaining stealthy. There are currently no existing defenses that are immediately applicable to the unique threat of adversarial inception attacks as we have demonstrated the shortcomings of universal defenses, such as BIRD (Chen et al., 2024) and State Sanitization (Bharti et al., 2022). Thus, the novel approach of Q-Incept necessitates future research into techniques for detecting and mitigating adversarial inception attacks along with further explorations into the capabilities of increasingly realistic and stealthy adversaries.
Acknowledgements
This research was developed with funding from the Defense Advanced Research Projects Agency (DARPA), under contract W912CG23C0031, and by NSF awards CNS-2312875, CNS-2331081, and FMitF 2319500. We thank Simona Boboila, Peter Chin, Lisa Oakley, and Aditya Vikram Singh for discussing various aspects of this project.
Impact Statement
In this paper we develop a new class of training-time, backdoor poisoning attacks against deep reinforcement learning agents. Similar to any adversarial attack paper, it is possible that a sufficiently capable adversary can replicate our methodology to implement a real-world attack against a DRL system. Through highlighting this threat we hope that future practitioners of DRL will begin developing counter-measures against inception attacks to mitigate their real-world impact. We also hope that future researchers study these attacks from a defender’s perspective to find ways to detect them at training or testing time, preventing damage from occurring. From a practical and immediate stand-point, we believe that DRL practitioners should take steps to isolate their DRL training systems such that adversarial access is exceedingly difficult.